OpenStudy (anonymous):

-(-(-5(-3+2)7)) simplify

OpenStudy (anonymous):

35

OpenStudy (anonymous):

-(-(-5(-3+2)7)) =-5(-3+2)7 = 5(3-2)7 =35

OpenStudy (anonymous):

thank you